OpenStudy (anonymous):

determine two values of b that will make 64xcubed+bx+16 a perfect square trinomial

OpenStudy (anonymous):

\[64x^3+b*x+16\]

OpenStudy (anonymous):

a perfect square trinomial is a square of a binomial: \[(a+b)^2\]

OpenStudy (anonymous):

if i assume the equation to be 64x^2 + b*x + 16 --> (a + b)^2 = a^2 + 2*a*b + b^2

OpenStudy (anonymous):

ok b^2 = 16, so \[\sqrt{16} = \pm\]

OpenStudy (anonymous):

+/- 4

OpenStudy (anonymous):

a^2 = 64x^2, so let's keep a=8x.....now for the 2*a*b portion. we have a and b, but each value can be positive or negative....so the value are +/- 4 and +/- 8.

OpenStudy (anonymous):

so 2*a*b

OpenStudy (anonymous):

from your equation: [bx] = 2*a*b

OpenStudy (anonymous):

so b can vary between 2*(+/-4)*(+/-8)